Can this be made into an equation.
Tom has 51 cents MORE than Jerry. On the way home both spend the SAME amount on candy.
When they get home Tom has 64 cents and Jerry has 13 cents. Solve for either, money each spent,
total money spent or the total Tom and Jerry had before buying candy.

Answer by stanbon(75887)

Tom has 51 cents MORE than Jerry.
Jerry had "x" cents.
Tom had "x+51" cents
--------------------------------
On the way home both spend the SAME amount on candy.
Let the amount each spent be "y":
-----
When they get home Tom has 64 cents and Jerry has 13 cents.
Equations:
x + 51- y = 64
x - y = 13
-------
Solve for either, money each spent,
total money spent or the total Tom and Jerry had before buying candy.
----
Rearrange the equations:
x - y = 13
x - y = 13
------
There are infinite number of solutions.
Example:
If x = 100 , y = 87
If x = 50, y = 37
etc.
======================
Cheers,
Stan H.